OpenStudy (anonymous):

Verify the identity. cot^2(x)/(cosx+1)=(1-sinx)/sinx

OpenStudy (anonymous):

I really stink at identity solving, and I think I need a walk through of this one. I can usually figure them out but this one caught me like a deer in headlights.

OpenStudy (anonymous):

Convert everything into sine and cosine

OpenStudy (anonymous):

\[ \frac{\frac{\cos^2 x}{\sin^2x}}{\cos x+1}=\frac{1-\sin x}{\sin x} \]

OpenStudy (anonymous):

Right, that's what my trig notes tell me. thank you for your response! Once setting up the identity as such, how do you go about simplifying each side?

OpenStudy (anonymous):

Get rid of fractions by multiplying the denominators.

OpenStudy (anonymous):

How so? This problem has more fractions than I know what to do with! I'm sorry for seeming a little slow to catch on, but I really appreciate the walk through!

OpenStudy (anonymous):

Multiply by \(\cos x+1\)

OpenStudy (anonymous):

Alright, cool! So on the left side, we're still left with \[\cos^{2}x/\sin^{2}x\]

OpenStudy (anonymous):

Sorry, math is really rough for me without a visual representation. I don't think I'm well-versed enough to apply worded explanations to equations without some sort of figure!

OpenStudy (anonymous):

\[ \frac{\cos^2 x}{\sin^2 x} = \frac{(1-\sin x)(\cos x+1)}{\sin x} \]

OpenStudy (anonymous):

At this point you want to multiply both sides by \(\sin^2 x\) now.

OpenStudy (anonymous):

Right; that way we can get rid of the denominators on both sides at once, correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Awesome! Are we done once we get rid of the denominators? It's mostly a conceptual issue for me.

OpenStudy (anonymous):

Hold on...

OpenStudy (anonymous):

Sorry, we can't do it the way I've been saying to do it. I made a mistake.

OpenStudy (anonymous):

We cant start with the equation, we have to start with one side and get to the other side.

OpenStudy (anonymous):

We have to start with \[ \frac{\cot ^2x}{\cos x+1} \]

OpenStudy (anonymous):

Oh, okay! How should we approach it, then? And also, feel free to put your work on here. This is all for my own personal notes; I just picked a random problem out of my text book to help me understand.

OpenStudy (anonymous):

Alright, I follow so far. Sounds simple enough.

OpenStudy (anonymous):

To get rid of the fraction, sort of... we do 'multiplication by the conjugate

OpenStudy (anonymous):

Now the conjugate of \(\cos x +1\) is \(\cos x -1\)

OpenStudy (anonymous):

So we multiply top and bottom by it.

OpenStudy (anonymous):

O kay. I vaguely recall such a method from class but it looks like I opted out of writing it down. Big mistake on my part, it would seem!

OpenStudy (anonymous):

Doing that gives us \[ \frac{\cot^2x(\cos x-1)}{\cos^2x -1} \]

OpenStudy (anonymous):

Now when ever we see \(\cos^2\) or \(\sin ^2\) and a \(1\) near by, we want to try using Pythagorean identity.

OpenStudy (anonymous):

I follow you thus far.

OpenStudy (anonymous):

\[ 1=\sin^2x+\cos^2x \\ 1-\sin^2x =\cos^2x \\ -\sin^2x =\cos^2x-1 \]

OpenStudy (anonymous):

Alright, I'm familiar with this identity; it's the first one we learn! How do I apply this to our previous fraction?

OpenStudy (anonymous):

We change the denominator \[ \frac{\cot^2x(\cos x-1)}{-\sin^2x} \]

OpenStudy (anonymous):

Ohhhh! I see now! -sin^{2}x is interchangeable with cos^{2}x-1!

OpenStudy (anonymous):

Now what? Are we finished with the left side of the equation?

OpenStudy (anonymous):

Now \[ \cot^2x=\frac{\cos^2x}{\sin^2x} \]

OpenStudy (anonymous):

Right, that's what we had before. What do we do with it?

OpenStudy (anonymous):

It's on top of -sin^{2}x, right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Alright, just making sure I'm with you - and I am! Go on, please.

OpenStudy (anonymous):

I'm wondering it this identity is even true.

OpenStudy (anonymous):

Are you sure you wrote it down right?

OpenStudy (anonymous):

It's an oddly written identity, that's for sure. I'll see if I can find it in my online textbook. One moment please.

OpenStudy (anonymous):

Here it is.

OpenStudy (anonymous):

It's \(\csc x\) not \(\cos x\). Big difference.

OpenStudy (anonymous):

Oh no, I'm sorry!!!!!!!!! It's been a long day, I feel awful!

OpenStudy (anonymous):

\[ \csc x = \frac 1{\sin x} \]

OpenStudy (anonymous):

Okay, I follow

OpenStudy (anonymous):

Sorry, I must go.